how many elecrons should be added or removed from a neutral body of mass 10mg so that it may remain hanging in air in an electric field of strength 100N/C directed upwards.

how  many elecrons  should  be  added  or removed from a  neutral  body  of mass 10mg so that it  may remain  hanging in air in  an electric field of strength 100N/C directed upwards.

Grade:12

1 Answers

Achleshwar Sankhla
6 years ago
In equilibrium F=qE =mgq=mg/E=(10×10^-3×10)/100q=0.001 CSince q=ne n=q/e =.001/1.6×10^-19 = 6.25 × 10^15